OpenStudy (anonymous):

The following function defines a recursive sequence: f(0) = -3 f(1) = 6 f(n) = -2•f(n -1) - f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -3, 6, -9, -12, … -3, 20, -95, 480, … -3, 6, -9, 12, … -3, -20, -95, -480, …

myininaya (myininaya):

try to find f(2) replace all n's with 2 and tell me what you get after doing only that step

OpenStudy (anonymous):

ok

OpenStudy (neer2890):

so in f(n) put 2 inplace of n i.e.,n=2 and match the 3rd term and so on.

myininaya (myininaya):

have you replaced all the n's with 2's in the recursive function: f(n)=-2*f(n-1)-f(n-2) ?

OpenStudy (anonymous):

I'm almost done @myininaya

OpenStudy (neer2890):

ok. f(2)=-2f(2-1)-f(2-2) =-2f(1)-f(0) =-2*6-(-3) =-12+3=-9

OpenStudy (anonymous):

how did you get 6 @neer2890

OpenStudy (neer2890):

f(1)=6 and f(0)=-3 it's given in question

OpenStudy (anonymous):

ok

OpenStudy (neer2890):

in the same way now find f(3) to find the exact answer.

OpenStudy (anonymous):

so the answer is either a or c

OpenStudy (neer2890):

yeah..so to find the correct answer you have to find f(3) by placing 3 inplace of n in f(n)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@neer2890 i got c

OpenStudy (neer2890):

dang... you're correct...good

OpenStudy (anonymous):

thanks :)

OpenStudy (neer2890):

anytime...:)